OpenStudy (anonymous):

The radius of the magnetic disk in a 3.5-inch diskette is 1.68 in. Find the linear speed of a point on the circumference of the disk if it is rotating at a speed of 350 rev/min. Answer in units of in/sec

OpenStudy (anonymous):

you need to convert rev/min to radians per second this is the angular velocity then you need to use the formula relating linear speed to angular velocity

OpenStudy (anonymous):

this is v=wr where w = angular velocity and r = radius

OpenStudy (anonymous):

We're looking at the speed of a point (let's pretend its a piece of dust) right on the edge of a spinning disk. Every revolution, the point will travel a distance pi*D, the circumference of the disk. Multiply this by revolutions per minute, and you get the speed of the dust particle. Don't forget to convert minutes to seconds v = pi*D*rpm = pi*(1.68 in*2)*350 rev/min * 1min/60second = 61.6 in/sec